OpenStudy (anonymous):

The terminal side of an angle theta in standard position coincides with the line y=1/2x in quadrant I. find sin theta to the nearest thousandth.

OpenStudy (phi):

is this \[y= \frac{1}{2}x \]

OpenStudy (anonymous):

Yes.

OpenStudy (phi):

I would pick a nice number for x (2 comes to mind) and plot the (x,y) pair now we have the same problem as the others

OpenStudy (phi):

notice that they say quadrant I (because the line extends into quadrant III , they have to tell us which way to go on the line)

OpenStudy (anonymous):

ok so. 

OpenStudy (phi):

yes, except it is hard to read (it looks like it's in code)

OpenStudy (anonymous):

So then it would 1/1/sqrt(5) so because it wants it in decimal form the answer would be .444?

OpenStudy (anonymous):

lol!

OpenStudy (phi):

how did you get 0.444?

OpenStudy (anonymous):

sorry I meant .447.

OpenStudy (phi):

sin A = 1/sqrt(5) (not 1/1/sqrt(5) ) but you must know that because you got 0.447 which is correct

OpenStudy (anonymous):

oh.. I did 1/1/sqrt(5) and then switched it to 1/sqrt(5) because of the reciprical of 1/sqrt(5).

OpenStudy (anonymous):

Does what I did make sense or did I do something and just happen to get the correct answer?

OpenStudy (phi):

first, you should use the idea that a reciprocal is the number flipped. (Less chance of making a mistake) second, in this problem they want the sin A so just use the definition of sin A sin A= opp/hyp = 1/sqrt(5) no flip involved

OpenStudy (anonymous):

Oh. I guess I was still using csc. So I thought that it was 1/sin A and I did the reciprocal but I am not sure why.. Now that you explained that.

OpenStudy (anonymous):

Ok. Thank you so much!

OpenStudy (phi):

I assume you now know that 1/1/sqrt(5) and then switched it to 1/sqrt(5) is not correct 1/1/sqrt(5) is sqrt(5) If you wanted the csc A and you know sin A= 1/sqrt(5) then flip to get sqrt(5)/1= sqrt(5)

OpenStudy (anonymous):

Yes. Now I understand what I did wrong even though I happened upon the correct answer.